OpenStudy (anonymous):

5/2- 3x= 6/7 -4x 3x +4x= 6/7 +-5/2 7x= 6/7*2 + -5/2*7 7x=12/14 + 10/14 7x/7= 22/4 x= 5 1/2 What did I do wrong?

OpenStudy (anonymous):

whoah... could you clarify what you did? Maybe skip a line for each new step

OpenStudy (anonymous):

5/2- 3x= 6/7 -4x 3x +4x= 6/7 +-5/2 7x= 6/7*2 + -5/2*7 7x=12/14 + 10/14 7x/7 = 22/4 x= 5 1/2 What did I do wrong?

OpenStudy (anonymous):

In your third line... You've got to multiply the top and bottom of the fraction by that number. Explanation: You can multiply things by 1 all day and still have it be the same number. When you multiply a fraction by a number to get the denominators the same , you don't want to entirely change the fraction- just make it look different. So you can multiply something by 7/7, and have it only make the fraction look different, because 7/7=1

OpenStudy (anonymous):

4x-3x = 6/7-5/2 x= ((6*2)-(5*7)/14 = (12-35)/14 x = -23/14

OpenStudy (anonymous):

thanks for clarifying. that's one of my biggest mistakes not multiplying by 1 and changing it to what i need.